Company INC is making baseball gloves. Last month, with 60 workers, they made 1000 gloves. If they hire 15 more workers this month, how many gloves can be made at the same rate? What would my equation look like to solve this?
Found 2 solutions by katie bennett, Alan3354:
Answer by katie bennett(2)

60 workers 15 workers
----------- = ------------
1000 gloves x =


i was taught that if you set up a porportion how it is wrote in the problem or question then...after setting up the problem you

cross multiply
for example :
15X1000 = 15000
the you would divide your answer by the remaining number which is 60

so you would divide
15000/60= 250 .

i hope you understand this and if you have any other questions just ask .


Answer by Alan3354(69443)

Company INC is making baseball gloves. Last month, with 60 workers, they made 1000 gloves. If they hire 15 more workers this month, how many gloves can be made at the same rate? What would my equation look like to solve this?
------------------------
60 workers made 1000 gloves
75 workers is 60*1.25
1.25 times as many workers will make 1.25 times as many gloves
1000*1.25 = 1250 gloves
---------------------------
You can set up a proportion
60:1000::75:g
60g = 75000
g = 75000/60
g = 1250 gloves
++++++++++++++++++++++
The other tutor calculated the output for 15 workers, not for 15 MORE workers.